# A Photorealistic Dataset and Vision-Based Algorithm for Anomaly Detection During Proximity Operations in Lunar Orbit

### v4

@@ Abstract
NASA’s forthcoming Lunar Gateway space station, which will be uncrewed most of the time, will need to operate with an unprecedented level of autonomy. One key challenge is enabling the Canadarm3, the Gateway’s external robotic system, to detect hazards in its environment using its onboard inspection cameras. This task is complicated by the extreme and variable lighting conditions in space. In this paper, we introduce the visual anomaly detection and localization task for the space domain and establish a benchmark based on a synthetic dataset called ALLO (Anomaly Localization in Lunar Orbit). We show that state-of-the-art visual anomaly detection methods often fail in the space domain, motivating the need for new approaches. To address this, we propose MRAD (Model Reference Anomaly Detection), a statistical algorithm that leverages the known pose of the Canadarm3 and a CAD model of the Gateway to generate reference images of the expected scene appearance. Anomalies are then identified as deviations from this model-generated reference. On the ALLO dataset, MRAD surpasses state-of-the-art anomaly detection algorithms, achieving an AP score of 62.1% at the pixel level and an AUROC score of 74.9% at the image level. Given the low tolerance for risk in space operations and the lack of domain-specific data, we emphasize the need for novel, robust, and accurate anomaly detection methods to handle the challenging visual conditions found in lunar orbit and beyond.
@@ I Introduction
To enhance robotic autonomy, vision-based anomaly detection methods that can autonomously detect collision hazards, such as loose tools or debris, are of particular interest. Anomaly detection and localization involve identifying and pinpointing the locations of unexpected content in images that deviate from a set of expected inputs. Anomaly detection in the space domain is particularly challenging because of complex lighting conditions that result from the black background coupled with harsh direct solar illumination. Furthermore, the varied camera viewpoints used during operations near the space station can cause anomalies to fall into the shadow of the station and blend in with the background, making detection substantially more difficult.
@@
Although reliable automated anomaly detection would be highly valuable, existing datasets and methods do not directly address the problem of anomaly detection in the space domain. Specifically, a significant challenge in space anomaly detection is the absence of comprehensive, labelled datasets necessary for robust evaluation. Moreover, our comparative analysis demonstrates that existing approaches significantly underperform and fail to generalize effectively to this domain. To address these gaps, we introduce ALLO, a large-scale open-source anomaly detection dataset tailored for robotic proximity operations in space. Using ALLO, we establish a new benchmark by evaluating existing anomaly detection algorithms. In addition, we present MRAD, a novel, statistical algorithm specifically designed for this application. MRAD leverages camera pose information and a known CAD model of the station to perform per-image anomaly detection, achieving state-of-the-art performance on the ALLO benchmark. In summary, our main contributions are as follows:
  • we introduce ALLO, a synthetic dataset for vision-based anomaly detection in the space domain, and release its open-source generation pipeline to facilitate extensions and future research;
  • we establish a new anomaly detection benchmark using ALLO, evaluating recent anomaly detection algorithms and highlighting their shortcomings for the space environment;
  • we introduce Model Reference Anomaly Detection (MRAD), a statistical algorithm specifically designed for the space domain, and demonstrate its superior performance over existing anomaly detection algorithms on ALLO through comprehensive evaluation and ablation studies.
The remainder of this paper is structured as follows. In Section II, we review existing anomaly detection datasets and algorithms. In Section III, we detail the ALLO dataset and its rendering pipeline. We then describe our proposed MRAD algorithm in Section IV and evaluate its performance against other state-of-the-art methods on the ALLO dataset in Section V. The ALLO dataset and all open-source code are available from the anonymous repository at https://anonymous.4open.science/r/ALLO-00A8.
@@ II Related Work
Although anomaly detection is an emerging topic in the space domain, there is extensive work on visual datasets for space applications and, separately, on industrial anomaly detection. In this section, we review existing image datasets for these two applications. In addition, we provide an overview of modern anomaly detection methods which will serve as baselines for our benchmark.
@@ II-A2 Industrial Defect Inspection
The majority of recent efforts in anomaly detection have focused on finding manufacturing defects during visual industrial inspection. The MVTec 2D anomaly detection dataset [8] is the most widely used benchmark for this application due to its pixel-level annotations and diverse range of objects and textures [9]. However, similar to other anomaly detection datasets such as BTech [10] and Kolector [11], MVTec images are all captured from the same perspective, under conistent lighting and with simple backgrounds. In contrast, the ALLO dataset offers a more comprehensive and diverse set of scenes, including multiple views of the ISS with varied lighting conditions and backgrounds.
@@ II-B2 Deep Learning-Based Detectors
Most of the modern deep learning-based anomaly detectors are unsupervised due to the scarcity of labelled anomalies, although some self-supervised approaches introduce synthetic anomalies during training [18, 19]. Unsupervised methods [20, 21, 22] learn a representation of normal data and detect deviations at inference time. Many operate in feature space, using pre-trained backbones (e.g., ResNet) and either fit statistical models to extracted features [20, 22] or estimate their density with normalizing flows [23, 24]. Alternatively, generative approaches reconstruct only normal images, flagging anomalies by their reconstruction error [9]. Student–teacher architectures [25, 21] follow this paradigm by distilling normal data representations from a teacher to a student, which fails to reproduce unseen anomalous patterns.
@@ III-A Scene Modelling
We use a high-fidelity Blender model of the ISS [26] as a stand-in for the Lunar Gateway, as no detailed Gateway model is currently available. Given the expected architectural and surface similarities of the Gateway to the ISS, this model provides a realistic approximation of the visual environment anticipated for the Canadarm3’s operations. The Gateway’s anticipated orbit [27] is approximated by an ellipse matching the expected perilune and apolune.
@@
To simulate the arm-mounted camera perspectives, 50 unique base poses are manually defined around the ISS. To introduce realistic pose imprecision and enhance viewpoint diversity, each base pose is perturbed with Gaussian noise during rendering, using standard deviations of 1 m for position and 0.2 rad for orientation (roll, pitch, and yaw). This randomization strategy ensures comprehensive coverage of the relevant station regions and subtle yet critical variations in perspective.
@@ III-B Image Rendering
Different camera positions are used for the training and test sets to avoid viewpoint overlap. The test set contains 16 distinct anomalous objects representing potential hazards, including dislodged station equipment or items lost during extravehicular activities (e.g., thermal blankets, cables, tools), as well as objects resembling small experimental satellites from prior missions (Fig. 2).
@@
Each rendered image has a resolution of 1,920 $\times$ 1,080 pixels and is paired with a three-class segmentation mask delineating the background, non-anomalous foreground (e.g., station surfaces, celestial bodies), and anomalous regions. The rendering pipeline, illustrated in Fig. 1, is repeated with two random seeds to expand the dataset. The dataset composition is reported in Table I, with sample images shown in Fig. 3.
@@ IV Model Reference Anomaly Detection
MRAD is an anomaly detection algorithm tailored for the space domain. It detects anomalies by comparing a query image with a reference image rendered from a known camera pose and a CAD model. The algorithm operates in three stages: (1) pixel-level anomaly scores are computed using an extended version of the Reed–Xiaoli detector (RXD) applied to query–reference pairs; (2) region growing is used to cluster pixels into candidate anomalous regions, producing a binary mask; and (3) an image is classified as anomalous if the total detected area exceeds a threshold. In our work, both reference and query images are synthetically generated—we simulate the domain gap between real and synthetic data by introducing random variations in lighting and pose when rendering the reference images.
@@ IV-A Reed-Xiaoli Detector (RXD) Anomaly Score
While effective for relatively homogeneous scenes, RXD struggles with non-uniform backgrounds, such as space imagery, where lighting and texture vary significantly. To address this limitation, we adapt principles from the probabilistic anomaly detector (PAD) [16], which compares likelihoods under both nominal and anomalous distributions. In MRAD, the nominal (anomaly-free) distribution is derived directly from the reference image, while the query image provides statistics for potentially anomalous regions. Specifically, the anomaly-free distribution statistics $(\boldsymbol{\mu}_{0},\boldsymbol{\Sigma}_{0})$ are computed from the reference image, and the query distribution statistics $(\boldsymbol{\mu}_{1},\boldsymbol{\Sigma}_{1})$ are computed from the query image. The MRAD anomaly score for a pixel $\mathbf{x}$ is then defined as
$\mathrm{MRAD}(\mathbf{x})=\mathrm{RXD}_{\mathrm{ref}}(\mathbf{x};\boldsymbol{\mu}_{0},\boldsymbol{\Sigma}0)-\mathrm{RXD}_{\mathrm{query}}(\mathbf{x};\boldsymbol{\mu}_{1},\boldsymbol{\Sigma}_{1}),$ (2)
@@
To account for spatial variability in lighting and texture (e.g., the contrast between the black background of space and the station’s surface), both reference and query images are divided into 4 $\times$ 4 grids. Mean and covariance statistics are computed per grid, yielding localized anomaly scores that better model region-specific variations. Scores are scaled in the range [0, 255] to produce a grayscale MRAD map, where pixels with values below a predefined threshold are classified as anomaly-free and set to zero, and remaining high-scoring pixels are retained for further analysis.
@@ IV-B Noise Removal
First, a preliminary check identifies grid cells that are likely to contain a large anomaly: if more than 90% of the scores in a cell exceed the minimum MRAD score, the entire cell is marked as anomalous. Next, a kernel-based method removes noise caused by random score variations. A sliding-window kernel (with a size $\sim$10% of the grid dimensions) is used to compute the local standard deviation of anomaly scores; regions with a standard deviation below an empirical threshold (about 3–5%) are considered to be indicative of uniform background noise and their scores are set to zero. Finally, a Gaussian blur is applied to the noise-filtered score map to smooth local fluctuations and merge fragmented anomalous regions, improving region contiguity for subsequent detection.
@@
ScoreImage = $\mathrm{Zeros}(\mathrm{SizeOf}(\text{QueryImage}))$
AnomalyMap = $\mathrm{Zeros}(\mathrm{SizeOf}(\text{QueryImage}))$
for GridCell in QueryImage do
  for Pixel in GridCell do
   Score = $\mathrm{MRAD}(\text{Pixel})$ $\triangleright$ See Eq. (2)
   if Score >> MinMRADScore then
     ScoreImage[Pixel] = Score
   end if
  end for
  StdDev = $\mathrm{StandardDeviation}(\text{GridCell},\text{Kernel})$
  if $\mathrm{Range}(\text{StdDev})<$ 3% then
   Set Score = 0 for all pixels in GridCell
  end if
end for
Apply Gaussian blur to ScoreImage
Apply mean shift clustering to ScoreImage
Seeds = $\mathrm{FindClusterCentroids}(\text{ScoreImage})$
for StartPoint in Seeds do
  AnomalyMap[StartPoint] = 1 $\triangleright$ Anomalous pixel
  for Pixel in $\mathrm{Footprint}(\text{StartPoint})$ do
   if $\mathrm{InBounds}(\text{Pixel, Tolerance})$ then
     AnomalyMap[Pixel] = 1 $\triangleright$ Anomalous pixel
   end if
  end for
end for
Count = $\mathrm{Sum}(\text{AnomalyMap})$
if Count >> AnomalyThreshold then
  ImageClass = Anomalous
else
  ImageClass = Normal
end if
@@ IV-C Clustering via Region Growing
Region growing then expands from each seed, aggregating spatially adjacent pixels with similar anomaly scores until no further candidates meet the criteria. This yields a binary anomaly map, from which the total number of anomalous pixels is computed. An image is classified as anomalous if this count exceeds a predefined threshold. The complete MRAD pipeline is outlined in Algorithm 1, with an illustrative example in Fig. 5.
@@
Figure 5: Visualization of the steps in the MRAD algorithm. The query image (a) and reference image (b) are the inputs to the algorithm. The anomaly score image (c), computed using Eq. 2, highlights the anomalous thermal blanket. Noise is removed and a Gaussian blur is applied, followed by mean-shift clustering to identify region growing seed points (d). Region growing aggregates anomalous pixels to generate the final anomaly map (e), accurately detecting the blanket while rejecting false positives. Notably, camera pose variations may introduce additional image differences that could be mistaken for anomalies if detection relied solely on pixel differences alone.
@@ V-A Experimental Setup
We modified the Intel Anomalib library [37] to load, train, and evaluate learning-based algorithms on the ALLO dataset. We selected eight representative learning-based methods from the library, chosen to cover both reconstruction- and representation-based approaches and for their established performance on the MVTec industrial benchmark.
Normalizing flow-based methods, FastFlow [24] and UFlow [38], learn the probability distribution of normal image features and detect anomalies as samples with low likelihood. Student–teacher methods, including STFPM [25] and Reverse Distillation [21], train a smaller student network to mimic the feature representations of a pre-trained teacher, flagging deviations as anomalies. Self-supervised and reconstruction-based methods, such as DRAEM [18], DSR [19], and SuperSimpleNet [39], reconstruct normal images or leverage synthetic anomalies to train models to discriminate between normal and abnormal data. Finally, CFA [22] is a feature-adaptation method that maps normal features to a shared hypersphere, detecting anomalies as features that fail to map correctly. Representation-based methods generally use fixed encoders pre-trained on ImageNet [40], while training remaining layers from scratch on the target dataset.
The proposed MRAD and the statistical detectors RXD and PAD were also evaluated on the ALLO test set with additional rendered reference images. RXD operates solely on the query image, while PAD and MRAD use both query and reference images; their anomaly scores were thresholded to produce output binary anomaly maps.
Ablation studies were also performed on MRAD to evaluate the effect of grid size and region-growing tolerance on detection performance. Grid size and tolerance were varied above and below their optimized values to assess robustness.
Performance was evaluated via image-level AUROC (I.AUROC), pixel-level AUROC (P.AUROC), and pixel-level average precision (P.AP). For pixel-level assessment on the imbalanced ALLO dataset, P.AUROC can be inflated by the predominant negative class, potentially obscuring localization failures. P.AP, which evaluates the precision-recall balance specifically for the anomalous class, offers a more robust metric for localization accuracy in this context [23]. Consequently, P.AP was the principal metric used for performance ranking.

### v5

@@ Abstract
NASA’s forthcoming Lunar Gateway space station, which will be uncrewed most of the time, will need to operate with an unprecedented level of autonomy. One key challenge is enabling the Canadarm3, the Gateway’s external robotic system, to detect hazards in its environment using its onboard inspection cameras. This task is complicated by the extreme and variable lighting conditions in space. In this paper, we introduce the visual anomaly detection and localization task for the space domain and establish a benchmark based on a synthetic dataset called ALLO (Anomaly Localization in Lunar Orbit). We show that state-of-the-art visual anomaly detection methods often fail in the space domain, motivating the need for new approaches. To address this, we propose MRAD (Model Reference Anomaly Detection), a statistical algorithm that leverages the known pose of the Canadarm3 and a CAD model of the Gateway to generate reference images of the expected scene appearance. Anomalies are then identified as deviations from this model-generated reference. On the ALLO dataset, MRAD surpasses state-of-the-art anomaly detection algorithms, achieving an AP score of 62.9% at the pixel level and an AUROC score of 75.0% at the image level. Given the low tolerance for risk in space operations and the lack of domain-specific data, we emphasize the need for novel, robust, and accurate anomaly detection methods to handle the challenging visual conditions found in lunar orbit and beyond.
@@ I Introduction
To enhance robotic autonomy, vision-based anomaly detection methods that can detect collision hazards, such as loose tools or debris, are of particular interest. Anomaly detection and localization refer to identifying and pinpointing image regions whose content deviates from the expected distribution of inputs. Anomaly detection in the space domain is particularly challenging because of complex lighting conditions that result from the black background coupled with harsh direct solar illumination. Furthermore, the varied camera viewpoints used during operations near the space station can cause anomalies to fall into shadow regions and blend in with the background, making detection substantially more difficult.
@@
Although reliable automated anomaly detection would be highly valuable, existing datasets and methods do not directly address the problem of anomaly detection in the space domain. Specifically, a significant challenge in space anomaly detection is the absence of comprehensive, labelled datasets necessary for robust evaluation. Moreover, our comparative analysis in this paper demonstrates that existing approaches significantly underperform, failing to generalize effectively to this domain. To address these gaps, we introduce ALLO, a large-scale open-source anomaly detection dataset tailored for robotic proximity operations in space. Using ALLO, we establish a benchmark by evaluating existing anomaly segmentation algorithms. We further present MRAD, a novel statistical algorithm specifically designed for this application that leverages camera pose information and a known CAD model of the station to perform per-image anomaly segmentation, achieving state-of-the-art performance on the ALLO benchmark. In summary, our main contributions are as follows.
  • We introduce ALLO, a synthetic dataset for vision-based anomaly detection in the space domain, and release its open-source generation pipeline to facilitate extensions and future research.
  • We establish an anomaly detection benchmark using ALLO, evaluating recent anomaly segmentation algorithms and highlighting their shortcomings for the space environment.
  • We introduce Model Reference Anomaly Detection (MRAD), a statistical algorithm designed for the space domain, and demonstrate its superior performance over existing anomaly detection algorithms on ALLO through comprehensive evaluation and ablation studies.
The remainder of the paper is structured as follows. In Section II, we review existing anomaly detection datasets and algorithms. In Section III, we detail the ALLO dataset and its rendering pipeline. We then describe our proposed MRAD algorithm in Section IV and evaluate its performance against other state-of-the-art methods on the ALLO dataset in Section V. The ALLO dataset and all open-source code are available at https://github.com/utiasSTARS/ALLO.
@@ II Related Work
Although anomaly detection is an emerging topic in the space domain, there is extensive work on visual datasets for space applications and, separately, on industrial anomaly detection. In this section, we review existing image datasets for these two applications. In addition, we provide an overview of modern anomaly detection methods that will serve as baselines for our benchmark.
@@ II-A2 Industrial Defect Inspection
The majority of recent efforts in anomaly detection have focused on finding manufacturing defects during visual industrial inspection. The MVTec 2D anomaly detection dataset [4] is the most widely used benchmark for this application due to its pixel-level annotations and diverse range of objects [40]. However, similar to other anomaly detection datasets such as BTech [24] and Kolektor [32], MVTec images are captured from fixed viewpoints, under consistent lighting and with simple backgrounds. In contrast, the ALLO dataset offers a more comprehensive and diverse set of scenes, including multiple views of the ISS with varied lighting conditions and complex backgrounds.
@@ II-B2 Deep Learning-Based Detectors
Most modern deep learning-based anomaly detectors are unsupervised due to the scarcity of labelled anomalies, although some self-supervised approaches introduce synthetic anomalies during training [38]. Unsupervised methods [9, 22] learn a representation of normal data and detect deviations at inference time. Many operate in feature space, using pre-trained backbones (e.g., ResNet) and either fit statistical models to extracted features [36, 22] or estimate their density with normalizing flows [13, 37]. In [31], an encoder-decoder framework is combined with anomaly scores similar to RXD to identify pixels far from a Gaussian distribution. Alternatively, generative approaches reconstruct only normal images, flagging anomalies by their reconstruction error [40]. Student–teacher architectures [35, 9] follow this paradigm by distilling normal data representations from a teacher to a student, thereby failing to reproduce unseen anomalous patterns.
@@ III-A Scene Modelling
We use a high-fidelity Blender model of the ISS [17] as a stand-in for the Lunar Gateway, since a detailed Gateway model is currently unavailable. Given the expected architectural and surface similarities of the Gateway to the ISS, this model provides a realistic approximation of the visual environment anticipated for the Canadarm3’s operations. The Gateway’s anticipated orbit [20] is approximated by an ellipse matching the expected perilune and apolune.
@@
To simulate arm-mounted camera perspectives, 50 unique base poses are manually defined around the ISS. These poses are derived from possible Canadarm3 trajectories near the mounting fixtures located on the exterior of the station. To introduce realistic pose imprecision and enhance viewpoint diversity, each base pose is perturbed with Gaussian noise during rendering, using standard deviations of 1 m for position and 0.2 rad for orientation (roll, pitch, and yaw). This randomization strategy ensures comprehensive coverage of the key station regions and subtle yet critical variations in perspective.
@@ III-B Image Rendering
Different camera positions are used for the training and test sets to avoid overlap between views. The test set contains 16 distinct anomalous objects representing potential collision hazards (see Fig. 2), including dislodged station equipment such as thermal blankets or cables, items lost during extravehicular activities such as tools or astronaut gloves, and objects resembling experimental equipment from prior missions such as Cubesats and small satellites.
@@
Each rendered image has a resolution of 1,920 $\mathbin{\!\!\times\!\!}$ 1,080 pixels and is paired with a three-class segmentation mask delineating the background, non-anomalous foreground (e.g., station surfaces, celestial bodies), and anomalous regions. The rendering pipeline, illustrated in Fig. 1, is repeated with two random seeds to expand the dataset. The dataset composition is reported in Table I, with sample images shown in Fig. 3.
@@ IV Model Reference Anomaly Detection
MRAD is an anomaly detection algorithm designed for the space domain. It detects anomalies by comparing a query image with a reference image rendered from a known camera pose and a CAD model of the station. The algorithm operates in three stages: (1) pixel-level anomaly scores are computed using an extended version of the Reed–Xiaoli detector (RXD) applied to query–reference pairs; (2) region growing clusters high-scoring pixels into candidate anomalous regions, producing a binary mask; and (3) an image is classified as anomalous if the total detected area exceeds a threshold. In this work, both reference and query images are synthetically generated, and the real-to-synthetic domain gap is simulated by introducing random variations in lighting and small perturbations to the camera pose when rendering the reference images.
@@ IV-A Reed-Xiaoli Detector (RXD) Anomaly Score
While effective for relatively homogeneous scenes, RXD struggles with non-uniform backgrounds, such as space imagery, where lighting and texture vary significantly. To address this limitation, we adapt principles from the probabilistic anomaly detector (PAD) [11], which compares likelihoods under both nominal and anomalous distributions. In MRAD, the nominal (anomaly-free) distribution is derived from the reference image, while the query image provides statistics for potentially anomalous regions. Specifically, the anomaly-free distribution statistics $(\boldsymbol{\mu}_{0},\boldsymbol{\Sigma}_{0})$ are computed from the reference image, and the query distribution statistics $(\boldsymbol{\mu}_{1},\boldsymbol{\Sigma}_{1})$ are computed from the query image. By leveraging statistics from both the reference and query images, MRAD can better handle complex, varying illumination and background structure, where the pixel distribution is non-Gaussian. The MRAD anomaly score for a pixel $\mathbf{x}$ is then defined as:
$\mathrm{MRAD}(\mathbf{x})=\mathrm{RXD}_{\mathrm{ref}}(\mathbf{x};\boldsymbol{\mu}_{0},\boldsymbol{\Sigma}_{0})-\mathrm{RXD}_{\mathrm{query}}(\mathbf{x};\boldsymbol{\mu}_{1},\boldsymbol{\Sigma}_{1}),$ (2)
@@
To account for spatial variability in lighting and texture (e.g., the contrast between the black background of space and the station’s surface), both reference and query images are divided into 4 $\mathbin{\!\!\times\!\!}$ 4 grids. Mean and covariance statistics are computed per grid, yielding localized anomaly scores that better model region-specific variations. Scores are scaled in the range [0, 255] to produce a grayscale MRAD map, where pixels with values below a predefined threshold are classified as anomaly-free and set to zero, and remaining high-scoring pixels are retained for further analysis. Fig. 5 illustrates a case where the pixel-intensity histogram for a reference grid cell is noticeably skewed and non-Gaussian, violating the Gaussian noise assumption. Nevertheless, in this instance MRAD remains robust, producing an accurate anomaly map that correctly localizes the free-floating cubesat.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 5: Left to right: reference image from the ALLO test set with one grid cell highlighted; pixel intensity histogram of the highlighted cell; anomalous query image; predicted anomaly map.
@@ IV-B Noise Removal
First, a preliminary check identifies grid cells that are likely to contain a large anomaly: if more than 90% of the scores in a cell exceed the minimum MRAD score, the entire cell is marked as anomalous. Next, a kernel-based method removes noise caused by random score variations. A sliding-window kernel (with a size $\sim$20% of the grid dimensions) is used to compute the local standard deviation of anomaly scores; regions with a standard deviation below an empirical threshold (about 3–5%) are considered to be indicative of uniform background noise and their scores are set to zero. Finally, a Gaussian blur is applied to the noise-filtered score map to smooth local fluctuations and merge fragmented anomalous regions, improving region contiguity for subsequent detection.
ScoreImage $\leftarrow$ Zeros(SizeOf(QueryImage));
AnomalyMap $\leftarrow$ Zeros(SizeOf(QueryImage));
foreach GridCell in QueryImage do
 foreach Pixel in GridCell do
    Score $\leftarrow$ MRAD(Pixel) ;
    // See Eq. (2)
    if Score >> MinMRADScore then
       ScoreImage[Pixel] $\leftarrow$ Score;
 StdDev $\leftarrow$ StdDeviation(GridCell, Kernel);
 if Range(StdDev) << 3% then
    Score = 0 for all pixels in GridCell;
Apply Gaussian blur to ScoreImage;
Apply mean shift clustering to ScoreImage;
Seeds $\leftarrow$ FindClusterCentroids(ScoreImage);
foreach StartPoint in Seeds do
 AnomalyMap[StartPoint] $\leftarrow$ 1 ;
 // Anomalous
 foreach Pixel in Footprint(StartPoint) do
    if InBounds(Pixel, Tolerance) then
       AnomalyMap[Pixel] $\leftarrow$ 1 ;
       // Anomalous
Count $\leftarrow$ Sum(AnomalyMap);
if Count >> AnomalyThreshold then
 ImageClass $\leftarrow$ Anomalous;
else
 ImageClass $\leftarrow$ Normal;
@@ IV-C Clustering via Region Growing
Region growing then expands from each seed, aggregating spatially adjacent pixels with similar anomaly scores until no further candidates meet the criteria. This yields a binary anomaly map, from which the total number of anomalous pixels is computed. An image is classified as anomalous if this count exceeds a predefined threshold. The complete MRAD pipeline is outlined in Algorithm 1 and Fig. 6.
@@
Figure 6: Visualization of the steps in the MRAD algorithm. The query image (a) and reference image (b) are the algorithm inputs. The anomaly score image (c), computed using Eq. 2, highlights the anomalous thermal blanket. Noise is removed and a Gaussian blur is applied, followed by mean-shift clustering to identify region growing seed points (d). Region growing aggregates anomalous pixels to generate the final anomaly map (e), detecting the blanket while rejecting false positives. Notably, camera pose variations may introduce additional image differences that could be mistaken for anomalies if detection relied solely on pixel differences alone.
@@ V-A Experimental Setup
We modified the Intel Anomalib library [3] to load, train, and evaluate learning-based algorithms on the ALLO dataset. We selected nine representative learning-based methods from the library, chosen to cover both reconstruction- and representation-based approaches and for their established performance on the MVTec industrial benchmark.
Normalizing flow-based methods, FastFlow [37] and UFlow [33], learn the probability distribution of normal image features and detect anomalies as samples with low likelihood. Student–teacher methods, including STFPM [35], Reverse Distillation [9] and UniNet [36] train a smaller student network to mimic the feature representations of a pre-trained teacher, flagging deviations as anomalies. UniNet introduces domain-aware feature selection, contrastive and margin losses, and a bottleneck with a weighted decision mechanism to support one model across multiple object categories without per-category training. Self-supervised and reconstruction-based methods, such as DSR [38], SuperSimpleNet [29], and Dinomaly [14], reconstruct normal images. Some approaches leverage synthetic anomalies for supervision, whereas Dinomaly uses a minimalist pure‑Transformer reconstructor. Finally, CFA [22] is a feature-adaptation method that maps normal features to a shared hypersphere, detecting anomalies as features that fail to map correctly. Representation-based methods generally use fixed encoders pre-trained on ImageNet [19] and train remaining layers from scratch on the target dataset.
The proposed MRAD algorithm and the statistical detectors RXD and PAD were also evaluated on the ALLO test set with additional rendered reference images. RXD operates solely on the query image, while PAD and MRAD use both query and reference images; their anomaly scores were thresholded to produce output binary anomaly maps. Ablation studies evaluated the impact of grid size, region-growing tolerance, kernel size, and the standard deviation used for noise removal on MRAD’s detection performance. Parameters were varied around their empirically optimized values to assess robustness.
Performance was measured using image-level AUROC (I.AUROC), pixel-level AUROC (P.AUROC), pixel-level average precision (P.AP), and the pixel-level false-positive rate at 95% true-positive rate (P.FPR95). On the imbalanced ALLO dataset, P.AUROC can be dominated by the negative class, obscuring localization failures. P.AP, which measures the precision-recall trade-off specifically for anomalous pixels, provides a more informative localization metric in this setting [13]. Consequently, P.AP was used as the primary metric for fine-grained anomaly segmentation, followed by P.FPR95.